### v1

# TennisTV: Do Multimodal Large Language Models Understand Tennis Rallies?

Abstract
Multimodal large language models (MLLMs) excel at general video understanding but struggle with fast, high-frequency sports like tennis, where rally clips are short yet information-dense. To systematically evaluate MLLMs in this challenging domain, we present TennisTV, the first and most comprehensive benchmark for tennis video understanding. TennisTV models each rally as a time-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation. It covers 8 tasks at rally and stroke levels and includes 2,500 human-verified questions. Evaluating 16 representative MLLMs, we provide the first systematic assessment of tennis video understanding. Results reveal substantial shortcomings and yield two key insights: (i) frame-sampling density should be tailored and balanced across tasks, and (ii) improving temporal grounding is essential for stronger reasoning.
Index Terms—  Multimodal Large Language Model, Sports Understanding, Video Analysis
1 Introduction
Tennis, renowned worldwide for its commercial impact and elite tournaments, has recently attracted significant research interest in applying artificial intelligence to tennis video understanding. However, despite the significant progress made in general video understanding[1, 2] and reasoning[3, 4], identifying fast-paced, high-frequency sports videos is still challenging[5]. As shown in the Fig 1, for human observers, even in fast-paced, high-frequency confrontations, they can directly depict the start and end of each hit, infer the type of hit and the trajectory of the ball. In contrast, even the most advanced MLLMs are struggling to deal with seemingly basic tasks, such as calculating the total number of hits in a rally.
In traditional video understanding, short videos typically require fewer frames due to their limited duration, whereas long videos generally rely on denser frame sampling. In contrast, tennis rally videos, though brief in duration, are highly information-dense. A single rally lasting fewer than 10 seconds may contain high-density, rapidly changing motion signals, such as a player’s stroke type choice, movement trajectory, the ball’s rebound position etc. The central challenge is that, despite their brevity, the frequent and high-speed dynamics of rallies demand dense frame sampling, often approaching the level required for long video analysis; otherwise, crucial hitting details are likely to be lost.
Fig. 1: Illustrations of a simple rally counting task easily handled by non-experts, but still challenging for the advanced commercial models.
In order to systematically evaluate the current MLLMs’ understanding of short, information-dense sports videos, as well as promote the community’s improvement of MLLMs’ understanding of tennis videos, we introduce TennisTV, the first high-quality benchmark specifically designed for understanding tennis rally videos. TennisTV defines a stroke event as the atomic unit, spanning movement initiation, stroke action choice, directional execution, and the ball’s landing. Each rally video is modeled as a time-ordered rally sequence composed of these consecutive stroke events. A total of 8 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements. We design an automated annotation pipeline. Through three-stage annotation process, we realize structured annotation integration for the entire rally based on the rough annotation provided by the original data source[5], extract additional athletes’ movement information, and generate questions. Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2,500 curated questions.
We conduct a comprehensive evaluation of 16 state of the art multimodal large language models (MLLMs) on TennisTV, including three closed-source commercial models, eight open-source non-thinking VideoLLMs, and five open-source thinking VideoLLMs. The study yields three findings: first, certain tasks exhibit a high degree of correlation; second, the optimal frame sampling density differs across tasks; third, explicit reasoning delivers gains on knowledge light tasks. From these insights above, we draw two practical takeaways: balance frame sampling density with task granularity, and to further enhance the temporal grounding capabilities of reasoning models. We hope that TennisTV and these insights will guide future advances in tennis video understanding.
Fig. 2: Overview and annotation pipeline of TennisTV Benchmark.
2 Related Work
2.1 Sports Understanding
Sports understanding [6] is a rapidly evolving field that encompasses multiple research topics and integrates diverse modalities, covering tasks such as action recognition [7], athlete analysis [8], tactical planning [9], foul recognition [10]. Recent work on MLLMs [11, 12, 13] has shifted toward general frameworks spanning diverse sports tasks, while existing work[5] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally-level context and causal relations. Consequently, MLLM capabilities in fast-paced, high-frequency sports video remain largely unexplored. This paper uses tennis as the empirical domain and proposes a novel benchmark for fast-paced, high-frequency sports video understanding, with the goal of advancing future research in this area.
2.2 MLLMs for Video
Inspired by the success of LLMs in natural language processing, researchers have begun to explore their extension to multimodal inputs, leading to the development of MLLMs. In image domain, [14, 15] have demonstrated that integrating visual encoders with LLMs can effectively process image-text pairs. However, video understanding presents greater challenges due to the temporal information[16]. Some early approaches extended LLMs to video domain by incorporating temporal modeling mechanisms[17]. Recent research has increasingly focused on improving the temporal and regional comprehension capabilities of Video MLLMs. With the emergence of Deepseek-R1[18], a growing body of work has begun to apply reinforcement learning (RL) to substantially enhance the reasoning abilities of MLLMs[3, 19, 4], thereby enabling stronger video understanding. However, little attention has been given to investigating how Video MLLMs interpret short, information-dense sports events. To address this gap, we conduct experiments using the carefully designed TennisTV to evaluate the benefits of enhancing base models for analyzing fast-pace, high-frequency sports videos.
3 TennisTV Benchmark
To rigorously evaluate the performance of models in tennis video analysis, we introduce Tennis Tour Video (TennisTV). The benchmark comprises 8 subtasks with 2500 questions, supporting multiple levels of video understanding. As shown in Fig 2, this section outlines benchmark overview and automatic annotation pipeline.
Table 1: Main results on TennisTV. AvgSE, AvgRS denotes the average metrics of tasks related to stroke event and rally sequence.
Method Frame Sampling AR MG HD HO AvgSE TP RC TI PW AvgRS Avg
Random - 25.0 20.0 20.0 33.3 24.6 25.0 25.0 25.0 50.0 31.3 27.9
[…]
Gemini2.5 Pro 1 FPS 47.4 41.7 34.0 45.0 42.0 48.6 31.5 45.6 51.7 44.4 43.2
GPT-4.1 1 FPS 59.7 23.3 32.5 52.7 42.1 52.0 38.8 47.8 53.0 47.9 45.0
GPT-4o 1 FPS 56.6 18.3 34.8 52.3 40.5 49.2 39.8 47.0 56.3 48.1 44.3
[…]
Video-LLaVA-7B 8 Frames 25.2 19.4 16.3 32.3 23.3 23.7 23.3 25.2 50.0 30.6 26.9
LLaVA-OV-7B 32 Frames 36.3 14.0 30.0 41.0 30.3 36.3 25.3 40.4 49.0 37.8 34.0
mPLUG-Owl3 128 Frames 42.8 16.1 33.8 46.0 34.7 35.7 18.3 39.3 50.0 35.8 35.3
Qwen2VL-7B 32 Frames 29.5 18.9 34.5 40.3 30.8 43.1 24.8 44.8 50.0 40.7 35.7
InternVid2.5 8B 128 Frames 49.5 16.7 17.8 43.0 31.8 42.2 19.0 42.6 50.3 38.5 35.1
Qwen2.5VL-3B 32 Frames 39.7 18.3 32.0 40.3 32.6 38.8 20.5 45.2 50.0 38.6 35.6
Qwen2.5VL-7B 32 Frames 48.6 28.3 34.5 45.7 39.3 41.9 22.3 45.6 48.0 39.5 39.4
MiMoVL-7B SFT 32 Frames 45.5 22.2 31.8 38.3 34.5 36.3 31.8 34.4 46.0 37.1 35.8
[…]
Video-R1 32 Frames 45.9 22.8 28.8 42.0 34.9 43.7 28.3 43.7 52.0 41.9 38.4
VCR-7B 32 Frames 45.5 26.7 35.5 23.7 32.9 34.2 33.3 43.7 47.7 39.7 36.3
VideoRFT 32 Frames 46.8 21.7 35.3 45.0 37.2 35.1 25.5 42.6 45.7 37.2 37.2
VersaVid-R1 32 Frames 52.3 21.7 36.8 44.7 38.9 38.2 27.0 47.8 50.0 40.8 39.8
MiMoVL-7B RL 32 Frames 48.0 21.1 33.3 42.7 36.3 40.9 28.8 42.6 57.3 42.4 39.3
3.1 Benchmark Task Analysis
When viewing a tennis rally, we observe a rapid, high-frequency exchange of tactics and technique. Each rally is initiated by one player, who accelerates into position, selects a stroke type, chooses a shot direction, and ultimately produces a specific landing point or outcome. After the ball crosses the net or rebounds, the opponent proceeds through the same process. The event sequence alternates between players until a stroke ends the point. Based on this understanding of the sport, TennisTV defines a stroke event as the atomic unit, spanning movement initiation, shot selection, directional execution, and ball landing. Each rally video is modeled as a time-ordered rally sequence composed of these consecutive stroke events.
Stroke Event. The stroke-level perspective targets fine-grained phenomena, where the model must recognize and localize specific stroke events such as Action Recognition (AR), Movement Grounding (MG), Hit Direction (HD) and Hit Outcome (HO). These tasks compel the model to detect subtle spatiotemporal cues at the frame level during rapid, interactive exchanges.
Rally Sequence. The rally-level perspective centers on the macro dynamics of a point and encompasses Technique Identification (TI), Rally Count (RC), Technique Preference (TP) and Point Winner determination (PW). These tasks require a model to analyze the entire game process and perceive each stroke event, so as to grasp the overall trend of the rally and the players’ technical tendencies.
Overall, TennisTV offers a comprehensive and challenge benchmark that probes both basic perception (e.g., AR, HD) and higher-level tactical reasoning (e.g., PW, TP), providing a robust, diversified testbed for video understanding in fast-paced, high-frequency sports.
3.2 Dataset Annotation Pipeline
TennisTV is derived from F3Set[5], which provides manual annotations of individual shots from tennis tournament videos. Based on these annotations, we perform a three-stage data cleaning and annotation pipeline.
Basic Data Preparation We first aggregated the shot annotations corresponding to the same rally in temporal order and generated basic rally-level technical statistics, such as rally length and the distribution of different strokes used by both players, to construct structured rally information. Next, we removed videos with an insufficient number of rallies from the original dataset. Moreover, because tennis tournaments are played on courts with diverse surfaces, the varying background colors can reduce viewing clarity. To address this issue, we employed Qwen2.5VL-7B[20] as a video quality judge model to filter out videos with poor visual quality.
Motion Information Extraction Secondly, the original annotations primarily capture the ball’s trajectory within individual stroke events, while overlooking the player’s movement direction and positional changes throughout the rally. To address this limitation, we developed a player motion extraction method to capture movement information between stroke events. For each video, we extracted $m$ originally annotated stroke frames from the structured rally information obtained in the first stage. We then paired adjacent frames to form m1m-1 stroke frame pairs. For each pair, we applied a lightweight object detector to localize the player’s position. By computing the $\Delta x$ and $\Delta y$ offsets between the bounding box coordinates across the two frames, we determined movement direction: if either offset exceeded a specific threshold $\delta$, the player is considered to have moved in that direction. Finally, the extracted player motion information was integrated into the structured rally information.
Question Generation We employed GPT-4o to aggregate the structured rally information and generated a comprehensive rally sequence description for each video. Using the rally descriptions and basic statistics, we generated 8 task questions using either GPT-4o or a predefined template. All questions were manually verified, resulting in a total of 2500 questions across 1298 videos.
4 Experiment
4.1 Implementation Details
We evaluate 13 open-source MLLMs that can process multi-frame video inputs, including 8 non-thinking models: Video-LLaVA [17] , LLaVA-OneVision-7B, Qwen2.5VL-3B&7B [20] , Qwen2VL-7B, mPLUG-Owl3 [21] , InternVideo2.5 [2] and MiMoVL-SFT-7B-2508 [22] , as well as 5 thinking models: Video-R1 [23] , VideoChat-R1 [3] , VideoRFT [4] , VersaVid-R1 [19] and MiMoVL-RL-7B-2508 [22] . In addition, we assess GPT-4o, GPT-4.1 and Gemini2.5-Pro through their APIs. For the open-source models, we followed the input frame settings in VideoMME [24] . For the closed-source models, we adopted a video sampling rate of 1 FPS for evaluation due to budget consideration.
4.2 Evaluation Metric
We adopt two different metrics to evaluate the model performance on TennisTV. For PW task, which aims to identify which player won the rally, the model is provided with a statement describing the outcome of the rally. The model’s answer is considered correct only if it correctly classifies the statement as True or False. For other tasks, we formulate them as multiple-choice questions and evaluate whether the model’s selected option is correct.
4.3 Main Results
Results Table 3 reports results on TennisTV. Among closed-source models, GPT-4.1 ranks first overall at 45.0%, edging GPT-4o and Gemini 2.5 Pro. Open-source systems are competitive: the strongest non-thinking baseline, Qwen2.5-VL-7B, reaches 39.4%, and the thinking variant VersaVid-R1 nudges this to 39.8%. Reasoning brings targeted gains on the same 7B backbone (RC improves from 22.3% to 27.0% and TI from 45.6% to 47.8%) while averages remain close. Leadership is task-specific: Gemini 2.5 Pro leads MG (41.7%), GPT-4.1 leads TP (52.0%), and GPT-4o attains the highest AvgRS (48.1%). Despite these gains, overall accuracies remain modest and only moderately above the random baseline, indicating substantial headroom for progress.
Fig. 3: Correlation matrix between 8 tasks in TennisTV.
Finding I — High degree of correlation among the AR, TI, and TP tasks. As shown in Fig 3, we observe a strong positive association among AR, TI, and TP. Performance improvements on any one of these tasks are consistently accompanied by gains on the other two, and this pattern holds across model families and frame-sampling protocols. Conceptually, the three tasks constitute a coherent capability cluster: identifying a specific action provides the primitives for technique identification, which in turn supports inferring tactic preferences. Practically, this correlation implies that progress on one task is predictive of progress on the others, and that ablations targeting low-level attribution or temporal stitching can transfer to tactic inference. Taken together, this coherence offers empirical support for the soundness of our benchmark task design.
Table 2: Ablation study of different frame sampling strategies on Qwen2.5VL-7B. We report the AvgSE, AvgRS and Avg metrics.
Frame Sampling AvgSE AvgRS Avg
1 FPS 37.5 39.2 38.3
16 Frames 38.1 39.8 39.0
32 Frames 39.3 39.5 39.4
64 Frames 38.7 38.5 38.6
Finding II — Optimal frame sampling density differs by task. As shown in Table 2, with Qwen2.5-VL-7B, overall performance peaks at 32 frames and further increasing the number of input frames yields diminishing or even negative returns, with the rally sequence metrics particularly affected. These patterns indicate that finer temporal resolution primarily benefits brief, high-frequency stroke cues, while rally-level reasoning saturates earlier and is more sensitive to redundancy.
Fig. 4: Visualization of performance on three TennisTV benchmark tasks, which require no prior knowledge of tennis, based on the same base model.
Finding III — Reasoning brings gains on knowledge-light tasks. To disentangle the effect of reasoning from tennis-specific priors, we analyze three knowledge-light tasks (MG, HD, RC) that are deliberately designed to minimize dependence on specialized domain knowledge (e.g., scoring rules). As shown in Fig. 4, three reasoning variants (VideoChat-R1, VideoRFT and VersaVid-R1) are derived from reinforcement learning post-training applied to the same Qwen2.5-VL-7B. With the backbone, tokenization scheme, and video interface held constant, the observed improvements over the non-reasoning baseline can be attributed to enhanced reasoning rather than architectural differences or domain-specific priors. In addition, we employ the chain-of-thought prompting method [25] to evaluate the original reasoning capability of the base model. Relative gains are most pronounced on RC and HD, where multi-step analysis and explicit temporal/trajectory reasoning are critical, and are steadier but smaller on MG. This pattern supports the view that explicit reasoning chiefly aids information aggregation, while offering more limited benefits on purely temporal-oriented tasks.
5 Takeaways
Based on our findings, we highlight two takeaways for future improvements in tennis video understanding:
  • Balancing frame sampling density across tasks. For tennis video understanding, the sampling frequency that yields the best performance differs between coarse global tasks and fine-grained tasks. Future work should pursue a principled balance between event granularity and sampling density, so that models capture long range context without losing the detailed temporal cues required for rapid event recognition.
  • Enhancing the temporal understanding ability of video reasoning models. The experimental results have proved that reasoning can improve rally-level understanding, but it has not benefited from fine-grained temporal grounding task. This reveals that current video reasoning methods still lack sufficient capabilities for temporal perception.
6 Conclusion
In this paper, we introduce TennisTV, the first benchmark for tennis video understanding specifically designed to evaluate the capabilities of MLLMs. TennisTV models each rally as a time-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation. Unified evaluations over diverse MLLMs show that reasoning and careful frame sampling help, yet models still struggle with the sport’s fast, information-dense dynamics. We expect TennisTV to spur advances in hierarchical temporal reasoning and robust sports video understanding.

### v2

@@
 To systematically evaluate MLLMs in this challenging domain, we present TennisTV, the first and most comprehensive benchmark for tennis video understanding.
-TennisTV models each rally as a time-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation.
-It covers 8 tasks at rally and stroke levels and includes 2,500 human-verified questions.
-Evaluating 16 representative MLLMs, we provide the first systematic assessment of tennis video understanding.
+TennisTV models each rally as a temporal-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation.
+It covers 9 tasks from the stroke level to the rally level and includes 2943 human-verified questions.
+Evaluating 17 representative MLLMs, we provide the first systematic assessment of tennis video understanding.
 Results reveal substantial shortcomings and yield two key insights: (i) frame-sampling density should be tailored and balanced across tasks, and (ii) improving temporal grounding is essential for stronger reasoning.
@@
 The central challenge is that, despite their brevity, the frequent and high-speed dynamics of rallies demand dense frame sampling, often approaching the level required for long video analysis; otherwise, crucial hitting details are likely to be lost.
-Fig. 1: Illustrations of a simple rally counting task easily handled by non-experts, but still challenging for the advanced commercial models.
+Fig. 1: Illustrations of a simple counting task easily handled by non-experts, but still challenging for the advanced commercial models.
 In order to systematically evaluate the current MLLMs’ understanding of short, information-dense sports videos, as well as promote the community’s improvement of MLLMs’ understanding of tennis videos, we introduce TennisTV, the first high-quality benchmark specifically designed for understanding tennis rally videos.
@@
 Each rally video is modeled as a time-ordered rally sequence composed of these consecutive stroke events.
-A total of 8 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements.
+A total of 9 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements.
 We design an automated annotation pipeline.
 Through three-stage annotation process, we realize structured annotation integration for the entire rally based on the rough annotation provided by the original data source[5], extract additional athletes’ movement information, and generate questions.
-Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2,500 curated questions.
+Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2943 curated questions.
 We conduct a comprehensive evaluation of 16 state of the art multimodal large language models (MLLMs) on TennisTV, including three closed-source commercial models, eight open-source non-thinking VideoLLMs, and five open-source thinking VideoLLMs.
@@
 Sports understanding [6] is a rapidly evolving field that encompasses multiple research topics and integrates diverse modalities, covering tasks such as action recognition [7], athlete analysis [8], tactical planning [9], foul recognition [10].
-Recent work on MLLMs [11, 12, 13] has shifted toward general frameworks spanning diverse sports tasks, while existing work[5] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally-level context and causal relations.
+Recent work on MLLMs [11, 12, 13] has shifted toward general frameworks spanning diverse sports tasks, while existing work[5] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally level context and causal relations.
 Consequently, MLLM capabilities in fast-paced, high-frequency sports video remain largely unexplored.
@@
 3 TennisTV Benchmark
-To rigorously evaluate the performance of models in tennis video analysis, we introduce Tennis Tour Video (TennisTV).
-The benchmark comprises 8 subtasks with 2500 questions, supporting multiple levels of video understanding.
-As shown in Fig 2, this section outlines benchmark overview and automatic annotation pipeline.
-Table 1: Main results on TennisTV. AvgSE, AvgRS denotes the average metrics of tasks related to stroke event and rally sequence.
-Method Frame Sampling AR MG HD HO AvgSE TP RC TI PW AvgRS Avg
-Random - 25.0 20.0 20.0 33.3 24.6 25.0 25.0 25.0 50.0 31.3 27.9
+To rigorously evaluate the performance of MLLMs in tennis video understanding, we introduce Tennis Tour Video (TennisTV).
+The benchmark comprises 9 subtasks with 2943 questions, supporting multiple levels of video understanding.
+As shown in Fig 2, this section outlines benchmark overview introduction and the automatic annotation pipeline.
+Table 1: Main results on TennisTV. AvgSE and AvgRS denote the average metrics of tasks related to stroke event and rally sequence.
+Method Frame Sampling MG AR HD HO AvgSE TP RC TI PW ErrA AvgRS Avg
+Random - 20.0 25.0 20.0 33.3 24.6 25.0 25.0 25.0 50.0 50.0 35.0 30.4
+Human-level - 47.8 100.0 96.3 93.3 84.4 91.6 92.5 98.0 94.3 87.5 92.8 89.0
 […]
-Gemini2.5 Pro 1 FPS 47.4 41.7 34.0 45.0 42.0 48.6 31.5 45.6 51.7 44.4 43.2
-GPT-4.1 1 FPS 59.7 23.3 32.5 52.7 42.1 52.0 38.8 47.8 53.0 47.9 45.0
-GPT-4o 1 FPS 56.6 18.3 34.8 52.3 40.5 49.2 39.8 47.0 56.3 48.1 44.3
+Gemini 2.5 Pro 1 FPS 41.7 47.4 34.0 45.0 42.0 49.1 31.5 51.0 51.7 54.6 47.6 45.1
+GPT-4.1 1 FPS 23.3 59.7 32.5 52.7 42.1 51.6 38.8 49.3 53.0 54.6 49.5 46.2
+GPT-4o 1 FPS 18.3 56.6 34.8 52.3 40.5 47.5 39.8 50.3 56.3 54.1 49.6 45.6
 […]
-Video-LLaVA-7B 8 Frames 25.2 19.4 16.3 32.3 23.3 23.7 23.3 25.2 50.0 30.6 26.9
-LLaVA-OV-7B 32 Frames 36.3 14.0 30.0 41.0 30.3 36.3 25.3 40.4 49.0 37.8 34.0
-mPLUG-Owl3 128 Frames 42.8 16.1 33.8 46.0 34.7 35.7 18.3 39.3 50.0 35.8 35.3
-Qwen2VL-7B 32 Frames 29.5 18.9 34.5 40.3 30.8 43.1 24.8 44.8 50.0 40.7 35.7
-InternVid2.5 8B 128 Frames 49.5 16.7 17.8 43.0 31.8 42.2 19.0 42.6 50.3 38.5 35.1
-Qwen2.5VL-3B 32 Frames 39.7 18.3 32.0 40.3 32.6 38.8 20.5 45.2 50.0 38.6 35.6
-Qwen2.5VL-7B 32 Frames 48.6 28.3 34.5 45.7 39.3 41.9 22.3 45.6 48.0 39.5 39.4
-MiMoVL-7B SFT 32 Frames 45.5 22.2 31.8 38.3 34.5 36.3 31.8 34.4 46.0 37.1 35.8
+Video-LLaVA-7B 8 Frames 19.4 25.2 16.3 32.3 23.3 24.8 23.3 25.0 50.0 50.0 34.6 29.6
+LLaVA-OneVision-7B 64 Frames 14.4 36.3 30.0 41.0 30.4 35.7 25.3 41.0 49.0 50.0 40.2 35.9
+mPLUG-Owl3 128 Frames 16.1 42.8 33.8 46.0 34.7 35.7 18.3 40.7 50.0 50.0 38.9 37.0
+Qwen2VL-7B 32 Frames 19.4 29.5 34.5 40.3 30.9 43.5 24.8 44.7 50.0 51.9 43.0 37.6
+InternVideo2.5-8B 128 Frames 20.0 49.5 17.8 43.0 32.6 42.5 19.0 43.3 50.3 49.3 40.9 37.2
+InternVL3-8B 64 Frames 31.9 34.8 31.3 45.3 35.8 40.7 30.3 44.3 50.0 50.0 43.1 39.8
+Qwen2.5VL-3B 32 Frames 17.8 39.7 32.0 40.3 32.5 39.8 20.5 46.3 50.0 50.0 41.3 37.4
+Qwen2.5VL-7B 32 Frames 26.7 48.6 34.5 45.7 38.9 42.2 22.3 42.7 48.0 51.0 41.2 40.2
+MiMo-VL-7B SFT-2508 32 Frames 22.8 45.5 31.8 38.3 34.6 44.7 31.8 41.0 46.0 52.6 43.2 39.4
 […]
-Video-R1 32 Frames 45.9 22.8 28.8 42.0 34.9 43.7 28.3 43.7 52.0 41.9 38.4
-VCR-7B 32 Frames 45.5 26.7 35.5 23.7 32.9 34.2 33.3 43.7 47.7 39.7 36.3
-VideoRFT 32 Frames 46.8 21.7 35.3 45.0 37.2 35.1 25.5 42.6 45.7 37.2 37.2
-VersaVid-R1 32 Frames 52.3 21.7 36.8 44.7 38.9 38.2 27.0 47.8 50.0 40.8 39.8
-MiMoVL-7B RL 32 Frames 48.0 21.1 33.3 42.7 36.3 40.9 28.8 42.6 57.3 42.4 39.3
+Video-R1 32 Frames 21.7 45.9 28.8 42.0 34.6 39.4 28.3 40.7 52.0 47.4 41.6 38.5
+VCR-Thinking-7B 32 Frames 28.3 45.5 35.5 23.7 33.3 38.5 33.3 44.0 47.7 50.0 42.7 38.5
+VideoRFT 32 Frames 28.3 46.8 35.3 45.0 38.9 37.3 25.5 45.3 45.7 49.8 40.7 39.9
+VersaVid-R1 32 Frames 24.4 52.3 36.8 44.7 39.6 37.9 27.0 46.7 50.0 51.2 42.6 41.2
+MiMo-VL-7B RL-2508 32 Frames 22.8 48.0 33.3 42.7 36.7 36.7 28.8 49.3 57.3 49.5 45.3 41.5
 3.1 Benchmark Task Analysis
@@
 Stroke Event.
-The stroke-level perspective targets fine-grained phenomena, where the model must recognize and localize specific stroke events such as Action Recognition (AR), Movement Grounding (MG), Hit Direction (HD) and Hit Outcome (HO).
+The stroke level perspective targets fine-grained phenomena, where the model must recognize and localize specific stroke events such as Action Recognition (AR), Movement Grounding (MG), Hit Direction (HD) and Hit Outcome (HO).
 These tasks compel the model to detect subtle spatiotemporal cues at the frame level during rapid, interactive exchanges.
 Rally Sequence.
-The rally-level perspective centers on the macro dynamics of a point and encompasses Technique Identification (TI), Rally Count (RC), Technique Preference (TP) and Point Winner determination (PW).
-These tasks require a model to analyze the entire game process and perceive each stroke event, so as to grasp the overall trend of the rally and the players’ technical tendencies.
-Overall, TennisTV offers a comprehensive and challenge benchmark that probes both basic perception (e.g., AR, HD) and higher-level tactical reasoning (e.g., PW, TP), providing a robust, diversified testbed for video understanding in fast-paced, high-frequency sports.
+The rally level perspective centers on the macro dynamics of a point and encompasses Technique Identification (TI), Rally Count (RC), Technique Preference (TP), Point Winner determination (PW) and Error Analysis (EA).
+These tasks require a model to analyze the entire game process, so as to grasp the overall trend of the rally and the players’ technical tendencies.
+Overall, TennisTV provides a comprehensive and challenging benchmark.
+From a general understanding perspective, it encompasses both the perception of basic object motion information (HD, MG) and the comprehensive summarization of video events (RC).
+Within the sports domain, it further includes tactical understanding (TP) and the win or loss judgment based on tennis rules (PW).
 3.2 Dataset Annotation Pipeline
-TennisTV is derived from F3Set[5], which provides manual annotations of individual shots from tennis tournament videos.
-Based on these annotations, we perform a three-stage data cleaning and annotation pipeline.
-Basic Data Preparation We first aggregated the shot annotations corresponding to the same rally in temporal order and generated basic rally-level technical statistics, such as rally length and the distribution of different strokes used by both players, to construct structured rally information.
+TennisTV is derived from F3Set[5], which contains manually annotated shots from tennis tournament videos. Based on raw annotations, we perform a three-stage data annotation pipeline.
+Basic Data Preparation We first aggregated the shot annotations corresponding to the same rally in temporal order and generated basic rally level technical statistics, such as rally length and the distribution of different strokes used by both players, to construct structured rally information.
 Next, we removed videos with an insufficient number of rallies from the original dataset.
@@
 Question Generation We employed GPT-4o to aggregate the structured rally information and generated a comprehensive rally sequence description for each video.
-Using the rally descriptions and basic statistics, we generated 8 task questions using either GPT-4o or a predefined template.
-All questions were manually verified, resulting in a total of 2500 questions across 1298 videos.
+Using the rally descriptions and basic statistics, we generated 9 task questions using either GPT-4o or a predefined template.
+All questions were manually verified, resulting in a total of 2943 questions across 1392 videos.
 4 Experiment
 4.1 Implementation Details
-We evaluate 13 open-source MLLMs that can process multi-frame video inputs, including 8 non-thinking models: Video-LLaVA [17] , LLaVA-OneVision-7B, Qwen2.5VL-3B&7B [20] , Qwen2VL-7B, mPLUG-Owl3 [21] , InternVideo2.5 [2] and MiMoVL-SFT-7B-2508 [22] , as well as 5 thinking models: Video-R1 [23] , VideoChat-R1 [3] , VideoRFT [4] , VersaVid-R1 [19] and MiMoVL-RL-7B-2508 [22] .
+We evaluate 14 open-source MLLMs that can process multi-frame video inputs, including 9 non-thinking models: Video-LLaVA [17] , LLaVA-OneVision-7B, Qwen2.5VL-3B&7B [20] , Qwen2VL-7B, mPLUG-Owl3 [21] , InternVideo2.5 [2] and MiMoVL-SFT-7B-2508 [22] , as well as 5 thinking models: Video-R1 [23] , VideoChat-R1 [3] , VideoRFT [4] , VersaVid-R1 [19] and MiMoVL-RL-7B-2508 [22] .
 In addition, we assess GPT-4o, GPT-4.1 and Gemini2.5-Pro through their APIs.
@@
 We adopt two different metrics to evaluate the model performance on TennisTV.
-For PW task, which aims to identify which player won the rally, the model is provided with a statement describing the outcome of the rally.
+For PW and EA, the model is provided with a statement describing the outcome of the rally.
 The model’s answer is considered correct only if it correctly classifies the statement as True or False.
@@
 Results Table 3 reports results on TennisTV.
-Among closed-source models, GPT-4.1 ranks first overall at 45.0%, edging GPT-4o and Gemini 2.5 Pro.
-Open-source systems are competitive: the strongest non-thinking baseline, Qwen2.5-VL-7B, reaches 39.4%, and the thinking variant VersaVid-R1 nudges this to 39.8%.
-Reasoning brings targeted gains on the same 7B backbone (RC improves from 22.3% to 27.0% and TI from 45.6% to 47.8%) while averages remain close.
-Leadership is task-specific: Gemini 2.5 Pro leads MG (41.7%), GPT-4.1 leads TP (52.0%), and GPT-4o attains the highest AvgRS (48.1%).
-Despite these gains, overall accuracies remain modest and only moderately above the random baseline, indicating substantial headroom for progress.
-Fig. 3: Correlation matrix between 8 tasks in TennisTV.
+Among closed-source models, GPT-4.1 rank first with an overall score of 46.2%, narrowly surpassing GPT-4o and Gemini 2.5 Pro.
+Open-source systems also demonstrate strong competitiveness: under the non-thinking setting, Qwen2.5-VL-7B achieved a score of 40.2%, while the best thinking model MiMo-VL-7B improved this to 41.5% (+1.3pp).
+On average, open-source reasoning models outperform their non-reasoning counterparts but still lagged considerably behind leading closed-source commercial systems.
+Despite these gains, overall accuracy remains well below human performance and, on some tasks, even falls beneath the random-chance baseline, indicating substantial headroom for progress.
+Fig. 3: Correlation matrix between 9 tasks in TennisTV.
 Finding I — High degree of correlation among the AR, TI, and TP tasks.
@@
 Frame Sampling AvgSE AvgRS Avg
-1 FPS 37.5 39.2 38.3
-16 Frames 38.1 39.8 39.0
-32 Frames 39.3 39.5 39.4
-64 Frames 38.7 38.5 38.6
-Finding II — Optimal frame sampling density differs by task.
-As shown in Table 2, with Qwen2.5-VL-7B, overall performance peaks at 32 frames and further increasing the number of input frames yields diminishing or even negative returns, with the rally sequence metrics particularly affected.
-These patterns indicate that finer temporal resolution primarily benefits brief, high-frequency stroke cues, while rally-level reasoning saturates earlier and is more sensitive to redundancy.
-Fig. 4: Visualization of performance on three TennisTV benchmark tasks, which require no prior knowledge of tennis, based on the same base model.
-Finding III — Reasoning brings gains on knowledge-light tasks.
-To disentangle the effect of reasoning from tennis-specific priors, we analyze three knowledge-light tasks (MG, HD, RC) that are deliberately designed to minimize dependence on specialized domain knowledge (e.g., scoring rules).
+1 FPS 37.9 41.3 39.8
+16 Frames 38.0 41.7 40.0
+32 Frames 38.9 41.2 40.2
+64 Frames 39.1 39.4 39.3
+Finding II — The optimal frame sampling density varies across perceptual levels.
+As we discussed above, In fast-paced, high-frequency sports videos, the number of frames required varies across different levels of understanding.
+As shown in Table 2, performance with Qwen2.5-VL-7B peaks at 32 frames, while adding more frames yields diminishing or even negative returns, particularly for rally-sequence metrics.
+These results suggest that finer temporal resolution is most beneficial for capturing brief, high-frequency stroke cues, whereas rally level reasoning saturates earlier and is more sensitive to redundancy.
+Fig. 4: Visualization of performance on three TennisTV tasks, which require no prior knowledge of tennis, based on the same base model.
+Finding III — Reasoning not always helps.
+To disentangle the effect of reasoning from tennis-specific priors, we analyze three knowledge-light tasks (MG, HD, RC) that minimize dependence on specialized rules.
+We analyze Qwen2.5-VL-7B, the zero-shot chain-of-thought (CoT) prompting method [25] and three RL post-trained reasoning variants (VideoChat-R1, VideoRFT, VersaVid-R1).
 As shown in Fig.
-4, three reasoning variants (VideoChat-R1, VideoRFT and VersaVid-R1) are derived from reinforcement learning post-training applied to the same Qwen2.5-VL-7B.
-With the backbone, tokenization scheme, and video interface held constant, the observed improvements over the non-reasoning baseline can be attributed to enhanced reasoning rather than architectural differences or domain-specific priors.
-In addition, we employ the chain-of-thought prompting method [25] to evaluate the original reasoning capability of the base model.
-Relative gains are most pronounced on RC and HD, where multi-step analysis and explicit temporal/trajectory reasoning are critical, and are steadier but smaller on MG.
-This pattern supports the view that explicit reasoning chiefly aids information aggregation, while offering more limited benefits on purely temporal-oriented tasks.
+4, reasoning variants exhibit consistent gains over the non-reasoning baseline on RC and HD, where multi-step aggregation and explicit trajectory reasoning are critical.
+However, on MG task, which focuses on temporal grounding, reasoning model do not show consistent improvement over the baseline.
+This suggests that existing RL strategies primarily enhance general content understanding, while temporal perception remains a key shortcoming.
 5 Takeaways
@@
   • Enhancing the temporal understanding ability of video reasoning models.
-The experimental results have proved that reasoning can improve rally-level understanding, but it has not benefited from fine-grained temporal grounding task.
-This reveals that current video reasoning methods still lack sufficient capabilities for temporal perception.
+The experimental results have proved that reasoning can improve rally level understanding, but it has not benefited from fine-grained temporal grounding task.
+This suggests that future RL-based post-training should emphasize temporal understanding and alignment mechanisms rather than relying solely on general video reasoning.
 6 Conclusion
-In this paper, we introduce TennisTV, the first benchmark for tennis video understanding specifically designed to evaluate the capabilities of MLLMs.
+In this paper, we present TennisTV, the first benchmark for tennis video understanding tailored to evaluate MLLMs.
 TennisTV models each rally as a time-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation.

### v3

@@
 TennisTV models each rally as a temporal-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation.
-It covers 9 tasks from the stroke level to the rally level and includes 2943 human-verified questions.
+It covers 8 tasks from the stroke level to the rally level and includes 2527 human-verified questions.
 Evaluating 17 representative MLLMs, we provide the first systematic assessment of tennis video understanding.
-Results reveal substantial shortcomings and yield two key insights: (i) frame-sampling density should be tailored and balanced across tasks, and (ii) improving temporal grounding is essential for stronger reasoning.
+Results yield two key insights: (i) frame-sampling density should be tailored and balanced across tasks, and (ii) improving temporal grounding is essential for stronger reasoning.
 Index Terms—  Multimodal Large Language Model, Sports Understanding, Video Analysis
@@
 However, despite the significant progress made in general video understanding[16, 20] and reasoning[11, 19], identifying fast-paced, high-frequency sports videos is still challenging[15].
-As shown in the Fig 1, for human observers, even in fast-paced, high-frequency confrontations, they can directly depict the start and end of each hit, infer the type of hit and the trajectory of the ball.
+As shown in the Fig 1, for human observers, even in dynamic scenes, they can directly depict the start and end of each stroke, infer the type of stroke and the trajectory of the ball.
 In contrast, even the most advanced MLLMs are struggling to deal with seemingly basic tasks, such as calculating the total number of hits in a rally.
@@
 Through three-stage annotation process, we realize structured annotation integration for the entire rally based on the rough annotation provided by the original data source[15], extract additional athletes’ movement information, and generate questions.
-Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2943 curated questions.
-We conduct a comprehensive evaluation of 16 state of the art multimodal large language models (MLLMs) on TennisTV, including three closed-source commercial models, eight open-source non-thinking VideoLLMs, and five open-source thinking VideoLLMs.
+Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2527 curated questions.
+We conduct a comprehensive evaluation of 17 state of the art multimodal large language models (MLLMs) on TennisTV, including three closed-source commercial models, eight open-source non-thinking VideoLLMs, and five open-source thinking VideoLLMs.
 The study yields three findings: first, certain tasks exhibit a high degree of correlation; second, the optimal frame sampling density differs across tasks; third, explicit reasoning delivers gains on knowledge light tasks.
@@
 Recent work on MLLMs [9, 22, 21] has shifted toward general frameworks spanning diverse sports tasks, while existing work[15] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally level context and causal relations.
-Consequently, MLLM capabilities in fast-paced, high-frequency sports video remain largely unexplored.
+Consequently, MLLM capabilities in tennis video remain largely unexplored.
 This paper uses tennis as the empirical domain and proposes a novel benchmark for fast-paced, high-frequency sports video understanding, with the goal of advancing future research in this area.
@@
 To address this gap, we conduct experiments using the carefully designed TennisTV to evaluate the benefits of enhancing base models for analyzing fast-pace, high-frequency sports videos.
+Table 1: Main results on TennisTV. AvgSE and AvgRS denote the average metrics of tasks related to stroke event and rally sequence.
+Method Frame Sampling MG AR HD HO AvgSE TP RC TI PW AvgRS Avg
+Random - 20.0 25.0 20.0 33.3 24.6 25.0 25.0 25.0 50.0 31.3 27.9
+Human-level - 47.8 100.0 96.3 93.3 84.4 91.6 92.5 98.0 94.3 94.1 89.2
+[…]
+Gemini 2.5 Pro 1 FPS 41.7 47.4 34.0 45.0 42.0 49.1 31.5 51.0 51.7 45.8 43.9
+GPT-4.1 1 FPS 23.3 59.7 32.5 52.7 42.1 51.6 38.8 49.3 53.0 48.2 45.1
+GPT-4o 1 FPS 18.3 56.6 34.8 52.3 40.5 47.5 39.8 50.3 56.3 48.5 44.5
+[…]
+Video-LLaVA-7B 8 Frames 19.4 25.2 16.3 32.3 23.3 24.8 23.3 25.0 50.0 30.8 27.0
+LLaVA-OV-7B 64 Frames 14.4 36.3 30.0 41.0 30.4 35.7 25.3 41.0 49.0 37.8 34.1
+mPLUG-Owl3 128 Frames 16.1 42.8 33.8 46.0 34.7 35.7 18.3 40.7 50.0 36.2 35.4
+Qwen2VL-7B 32 Frames 19.4 29.5 34.5 40.3 30.9 43.5 24.8 44.7 50.0 40.8 35.8
+InternVideo2.5-8B 128 Frames 20.0 49.5 17.8 43.0 32.6 42.5 19.0 43.3 50.3 38.8 35.7
+InternVL3-8B 64 Frames 31.9 34.8 31.3 45.3 35.8 40.7 30.3 44.3 50.0 41.3 38.5
+Qwen2.5VL-3B 32 Frames 17.8 39.7 32.0 40.3 32.5 39.8 20.5 46.3 50.0 39.2 35.8
+Qwen2.5VL-7B 32 Frames 26.7 48.6 34.5 45.7 38.9 42.2 22.3 42.7 48.0 38.8 38.8
+MiMoVL-7B SFT 32 Frames 22.8 45.5 31.8 38.3 34.6 44.7 31.8 41.0 46.0 40.9 37.4
+[…]
+Video-R1 32 Frames 21.7 45.9 28.8 42.0 34.6 39.4 28.3 40.7 52.0 40.1 36.8
+VCR-Thinking-7B 32 Frames 28.3 45.5 35.5 23.7 33.3 38.5 33.3 44.0 47.7 40.9 38.9
+VideoRFT 32 Frames 28.3 46.8 35.3 45.0 38.9 37.3 25.5 45.3 45.7 38.5 38.1
+VersaVid-R1 32 Frames 24.4 52.3 36.8 44.7 39.6 37.9 27.0 46.7 50.0 40.4 40.6
+MiMoVL-7B RL 32 Frames 22.8 48.0 33.3 42.7 36.7 36.7 28.8 49.3 57.3 44.3 40.5
 3 TennisTV Benchmark
 To rigorously evaluate the performance of MLLMs in tennis video understanding, we introduce Tennis Tour Video (TennisTV).
-The benchmark comprises 9 subtasks with 2943 questions, supporting multiple levels of video understanding.
-As shown in Fig 2, this section outlines benchmark overview introduction and the automatic annotation pipeline.
-Table 1: Main results on TennisTV. AvgSE and AvgRS denote the average metrics of tasks related to stroke event and rally sequence.
-Method Frame Sampling MG AR HD HO AvgSE TP RC TI PW ErrA AvgRS Avg
-Random - 20.0 25.0 20.0 33.3 24.6 25.0 25.0 25.0 50.0 50.0 35.0 30.4
-Human-level - 47.8 100.0 96.3 93.3 84.4 91.6 92.5 98.0 94.3 87.5 92.8 89.0
-[…]
-Gemini 2.5 Pro 1 FPS 41.7 47.4 34.0 45.0 42.0 49.1 31.5 51.0 51.7 54.6 47.6 45.1
-GPT-4.1 1 FPS 23.3 59.7 32.5 52.7 42.1 51.6 38.8 49.3 53.0 54.6 49.5 46.2
-GPT-4o 1 FPS 18.3 56.6 34.8 52.3 40.5 47.5 39.8 50.3 56.3 54.1 49.6 45.6
-[…]
-Video-LLaVA-7B 8 Frames 19.4 25.2 16.3 32.3 23.3 24.8 23.3 25.0 50.0 50.0 34.6 29.6
-LLaVA-OneVision-7B 64 Frames 14.4 36.3 30.0 41.0 30.4 35.7 25.3 41.0 49.0 50.0 40.2 35.9
-mPLUG-Owl3 128 Frames 16.1 42.8 33.8 46.0 34.7 35.7 18.3 40.7 50.0 50.0 38.9 37.0
-Qwen2VL-7B 32 Frames 19.4 29.5 34.5 40.3 30.9 43.5 24.8 44.7 50.0 51.9 43.0 37.6
-InternVideo2.5-8B 128 Frames 20.0 49.5 17.8 43.0 32.6 42.5 19.0 43.3 50.3 49.3 40.9 37.2
-InternVL3-8B 64 Frames 31.9 34.8 31.3 45.3 35.8 40.7 30.3 44.3 50.0 50.0 43.1 39.8
-Qwen2.5VL-3B 32 Frames 17.8 39.7 32.0 40.3 32.5 39.8 20.5 46.3 50.0 50.0 41.3 37.4
-Qwen2.5VL-7B 32 Frames 26.7 48.6 34.5 45.7 38.9 42.2 22.3 42.7 48.0 51.0 41.2 40.2
-MiMo-VL-7B SFT-2508 32 Frames 22.8 45.5 31.8 38.3 34.6 44.7 31.8 41.0 46.0 52.6 43.2 39.4
-[…]
-Video-R1 32 Frames 21.7 45.9 28.8 42.0 34.6 39.4 28.3 40.7 52.0 47.4 41.6 38.5
-VCR-Thinking-7B 32 Frames 28.3 45.5 35.5 23.7 33.3 38.5 33.3 44.0 47.7 50.0 42.7 38.5
-VideoRFT 32 Frames 28.3 46.8 35.3 45.0 38.9 37.3 25.5 45.3 45.7 49.8 40.7 39.9
-VersaVid-R1 32 Frames 24.4 52.3 36.8 44.7 39.6 37.9 27.0 46.7 50.0 51.2 42.6 41.2
-MiMo-VL-7B RL-2508 32 Frames 22.8 48.0 33.3 42.7 36.7 36.7 28.8 49.3 57.3 49.5 45.3 41.5
+The benchmark comprises 8 subtasks with 2527 questions, supporting multiple levels of video understanding.
+This section outlines benchmark introduction and automatic annotation pipeline.
+Fig. 3: Benchmark statistics of TennisTV.
 3.1 Benchmark Task Analysis
@@
 Stroke Event.
-The stroke level perspective targets fine-grained phenomena, where the model must recognize and localize specific stroke events such as Action Recognition (AR), Movement Grounding (MG), Hit Direction (HD) and Hit Outcome (HO).
-These tasks compel the model to detect subtle spatiotemporal cues at the frame level during rapid, interactive exchanges.
+The stroke-level perspective addresses fine-grained details.
+Action Recognition (AR) requires the model to classify basic stroke types, while Movement Grounding (MG) targets the precise temporal localization of player movements.
+Moreover, Hit Direction (HD) emphasizes on the ball’s trajectory, and Hit Outcome (HO) asks the model to determine the immediate result of one stroke.
+These tasks jointly depend on the model’s ability to capture subtle spatiotemporal cues at the frame level.
 Rally Sequence.
-The rally level perspective centers on the macro dynamics of a point and encompasses Technique Identification (TI), Rally Count (RC), Technique Preference (TP), Point Winner determination (PW) and Error Analysis (EA).
-These tasks require a model to analyze the entire game process, so as to grasp the overall trend of the rally and the players’ technical tendencies.
-Overall, TennisTV provides a comprehensive and challenging benchmark.
-From a general understanding perspective, it encompasses both the perception of basic object motion information (HD, MG) and the comprehensive summarization of video events (RC).
-Within the sports domain, it further includes tactical understanding (TP) and the win or loss judgment based on tennis rules (PW).
+The rally-level perspective captures the macro dynamics of a rally.
+Technique Identification (TI) requires the model to detect specific tactics, while Rally Count (RC) tasks it with recording the total number of strokes.
+Technique Preference (TP) demands that the model analyze in-rally tactical choices to reveal player tendencies, and Point Winner (PW) requires the model to identify the winner of the point.
+These tasks collectively necessitate a holistic analysis to capture overarching technical patterns.
+Overall, TennisTV offers a comprehensive benchmark that probes both basic perception (e.g., AR, HD) and tactical reasoning (e.g., PW, TP), providing a robust, diversified testbed for video understanding in fast-paced, high-frequency sports.
 3.2 Dataset Annotation Pipeline
-TennisTV is derived from F3Set[5], which contains manually annotated shots from tennis tournament videos. Based on raw annotations, we perform a three-stage data annotation pipeline.
-Basic Data Preparation We first aggregated the shot annotations corresponding to the same rally in temporal order and generated basic rally level technical statistics, such as rally length and the distribution of different strokes used by both players, to construct structured rally information.
+TennisTV is derived from F3Set[15], which provides manual annotations of individual shots from tennis tournament videos. Based on these annotations, we perform a three-stage annotation pipeline.
+Basic Data Preparation We first aggregated the shot annotations corresponding to the same rally in temporal order and generated basic rally-level technical statistics, such as rally length and the distribution of different strokes used by both players, to construct structured rally information.
 Next, we removed videos with an insufficient number of rallies from the original dataset.
@@
 Question Generation We employed GPT-4o to aggregate the structured rally information and generated a comprehensive rally sequence description for each video.
-Using the rally descriptions and basic statistics, we generated 9 task questions using either GPT-4o or a predefined template.
-All questions were manually verified, resulting in a total of 2943 questions across 1392 videos.
+Using the rally descriptions and basic statistics, we generated 8 task questions using either GPT-4o or a predefined template.
+All questions were manually verified, resulting in a total of 2527 questions across 1298 videos.
 4 Experiment
 4.1 Implementation Details
-We evaluate 14 open-source MLLMs that can process multi-frame video inputs, including 9 non-thinking models: Video-LLaVA [17] , LLaVA-OneVision-7B, Qwen2.5VL-3B&7B [20] , Qwen2VL-7B, mPLUG-Owl3 [21] , InternVideo2.5 [2] and MiMoVL-SFT-7B-2508 [22] , as well as 5 thinking models: Video-R1 [23] , VideoChat-R1 [3] , VideoRFT [4] , VersaVid-R1 [19] and MiMoVL-RL-7B-2508 [22] .
-In addition, we assess GPT-4o, GPT-4.1 and Gemini2.5-Pro through their APIs.
-For the open-source models, we followed the input frame settings in VideoMME [24] .
-For the closed-source models, we adopted a video sampling rate of 1 FPS for evaluation due to budget consideration.
+We evaluate 14 open-source MLLMs that supports multi-frame inputs, including 9 non-thinking models: Video-LLaVA [12], LLaVA-OV-7B, Qwen2VL-7B, Qwen2.5VL-3B&7B [1], mPLUG-Owl3 [24], InternVideo2.5 [20], InternVL3-8B and MiMoVL-SFT-7B-2508, as well as 5 thinking models: Video-R1 [5] ,VideoChat-R1 [11] ,VideoRFT [19] ,VersaVid-R1 [2] and MiMoVL-RL-7B-2508 [23].
+We assess GPT-4o, GPT-4.1, Gemini2.5-Pro through their APIs.
+For open-source models, we followed the input frame settings in VideoMME [6].
+For closed-source models, we adopted 1 FPS video sampling rate for evaluation due to budget consideration.
 4.2 Evaluation Metric
-We adopt two different metrics to evaluate the model performance on TennisTV.
-For PW and EA, the model is provided with a statement describing the outcome of the rally.
-The model’s answer is considered correct only if it correctly classifies the statement as True or False.
-For other tasks, we formulate them as multiple-choice questions and evaluate whether the model’s selected option is correct.
+We assess performance on TennisTV using two formats.
+In PW task, the model must correctly classify a statement about the rally outcome as True or False.
+All remaining tasks are cast as multiple-choice questions, evaluated based on the selection of the correct answer.
 4.3 Main Results
-Results Table 3 reports results on TennisTV.
-Among closed-source models, GPT-4.1 rank first with an overall score of 46.2%, narrowly surpassing GPT-4o and Gemini 2.5 Pro.
-Open-source systems also demonstrate strong competitiveness: under the non-thinking setting, Qwen2.5-VL-7B achieved a score of 40.2%, while the best thinking model MiMo-VL-7B improved this to 41.5% (+1.3pp).
-On average, open-source reasoning models outperform their non-reasoning counterparts but still lagged considerably behind leading closed-source commercial systems.
-Despite these gains, overall accuracy remains well below human performance and, on some tasks, even falls beneath the random-chance baseline, indicating substantial headroom for progress.
-Fig. 3: Correlation matrix between 9 tasks in TennisTV.
+Results Table 2.2 shows results on TennisTV.
+Among closed-source models, GPT-4.1 rank first with an overall score of 45.1%, narrowly surpassing GPT-4o and Gemini 2.5 Pro.
+Open-source systems also demonstrate strong competitiveness: Qwen2.5-VL-7B achieves a score of 38.8%, while the best thinking model VersaVid-R1 improves this to 40.6% (+1.8pp).
+On average, open-source reasoning models outperform their non-reasoning counterparts but still lag considerably behind leading closed-source commercial systems.
+Despite these gains, overall accuracy remains well below human performance and even falls beneath the random baseline on some tasks, indicating substantial headroom for progress.
+Fig. 4: Correlation matrix between 8 tasks in TennisTV.
 Finding I — High degree of correlation among the AR, TI, and TP tasks.
-As shown in Fig 3, we observe a strong positive association among AR, TI, and TP.
-Performance improvements on any one of these tasks are consistently accompanied by gains on the other two, and this pattern holds across model families and frame-sampling protocols.
-Conceptually, the three tasks constitute a coherent capability cluster: identifying a specific action provides the primitives for technique identification, which in turn supports inferring tactic preferences.
-Practically, this correlation implies that progress on one task is predictive of progress on the others, and that ablations targeting low-level attribution or temporal stitching can transfer to tactic inference.
-Taken together, this coherence offers empirical support for the soundness of our benchmark task design.
+Fig 4 reveals a robust positive correlation among AR, TI, and TP that persists across diverse model families and sampling protocols.
+We attribute this to the fact that these tasks form a coherent capability hierarchy: recognizing specific actions yields the primitives for technique identification, which subsequently supports the inference of tactic preferences.
+Consequently, performance on one task is highly predictive of the others, suggesting that low-level refinements in attribution or temporal stitching can transfer to tactic inference.
+Taken together, these findings underscore the empirical soundness of our benchmark.
 Table 2: Ablation study of different frame sampling strategies on Qwen2.5VL-7B. We report the AvgSE, AvgRS and Avg metrics.
 Frame Sampling AvgSE AvgRS Avg
-1 FPS 37.9 41.3 39.8
-16 Frames 38.0 41.7 40.0
-32 Frames 38.9 41.2 40.2
-64 Frames 39.1 39.4 39.3
+1 FPS 37.9 38.3 38.1
+16 Frames 38.0 39.3 38.6
+32 Frames 38.9 38.8 38.8
+64 Frames 39.1 37.7 38.4
 Finding II — The optimal frame sampling density varies across perceptual levels.
-As we discussed above, In fast-paced, high-frequency sports videos, the number of frames required varies across different levels of understanding.
+In fast-paced, high-frequency sports videos, the number of frames required varies across different levels of understanding.
 As shown in Table 2, performance with Qwen2.5-VL-7B peaks at 32 frames, while adding more frames yields diminishing or even negative returns, particularly for rally-sequence metrics.
 These results suggest that finer temporal resolution is most beneficial for capturing brief, high-frequency stroke cues, whereas rally level reasoning saturates earlier and is more sensitive to redundancy.
-Fig. 4: Visualization of performance on three TennisTV tasks, which require no prior knowledge of tennis, based on the same base model.
+Fig. 5: Visualization of performance on three sub-tasks, which require no prior knowledge of tennis, based on the same base model.
 Finding III — Reasoning not always helps.
@@
 As shown in Fig.
-4, reasoning variants exhibit consistent gains over the non-reasoning baseline on RC and HD, where multi-step aggregation and explicit trajectory reasoning are critical.
+5, reasoning variants exhibit consistent gains over the non-reasoning baseline on RC and HD, where multi-step aggregation and explicit trajectory reasoning are critical.
 However, on MG task, which focuses on temporal grounding, reasoning model do not show consistent improvement over the baseline.
@@
 The experimental results have proved that reasoning can improve rally level understanding, but it has not benefited from fine-grained temporal grounding task.
-This suggests that future RL-based post-training should emphasize temporal understanding and alignment mechanisms rather than relying solely on general video reasoning.
+This suggests that future RL post-training should emphasize temporal understanding and alignment mechanisms rather than relying solely on general video reasoning.
 6 Conclusion
-In this paper, we present TennisTV, the first benchmark for tennis video understanding tailored to evaluate MLLMs.
-TennisTV models each rally as a time-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation.
-Unified evaluations over diverse MLLMs show that reasoning and careful frame sampling help, yet models still struggle with the sport’s fast, information-dense dynamics.
-We expect TennisTV to spur advances in hierarchical temporal reasoning and robust sports video understanding.
+In this paper, we present TennisTV, the first benchmark ever for tennis video understanding tailored to evaluate MLLMs.
+TennisTV models each rally as a time-ordered sequence of consecutive strokes, using automated pipelines for question generation.
+Evaluations over diverse MLLMs show that reasoning and careful frame sampling help, yet models still struggle with the sport’s fast, information-dense dynamics.
+We expect TennisTV to spur advances in dynamic sports video understanding.

### v4

@@
 TennisTV models each rally as a temporal-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation.
-It covers 8 tasks from the stroke level to the rally level and includes 2527 human-verified questions.
+It covers 8 tasks from the stroke-level to the rally-level and includes 2527 human-verified questions.
 Evaluating 17 representative MLLMs, we provide the first systematic assessment of tennis video understanding.
@@
 Each rally video is modeled as a time-ordered rally sequence composed of these consecutive stroke events.
-A total of 9 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements.
+A total of 8 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements.
 We design an automated annotation pipeline.
@@
 Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2527 curated questions.
-We conduct a comprehensive evaluation of 17 state of the art multimodal large language models (MLLMs) on TennisTV, including three closed-source commercial models, eight open-source non-thinking VideoLLMs, and five open-source thinking VideoLLMs.
+We conduct a comprehensive evaluation of 17 state of the art MLLMs on TennisTV, including three closed-source commercial models, eight open-source non-thinking VideoLLMs, and five open-source thinking VideoLLMs.
 The study yields three findings: first, certain tasks exhibit a high degree of correlation; second, the optimal frame sampling density differs across tasks; third, explicit reasoning delivers gains on knowledge light tasks.
@@
 Sports understanding [18] is a rapidly evolving field that encompasses multiple research topics and integrates diverse modalities, covering tasks such as action recognition [3], athlete analysis [17], tactical planning [14], foul recognition [4].
-Recent work on MLLMs [9, 22, 21] has shifted toward general frameworks spanning diverse sports tasks, while existing work[15] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally level context and causal relations.
+Recent work on MLLMs [9, 22, 21] has shifted toward general frameworks spanning diverse sports tasks, while existing work[15] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally-level context and causal relations.
 Consequently, MLLM capabilities in tennis video remain largely unexplored.
@@
 Consequently, performance on one task is highly predictive of the others, suggesting that low-level refinements in attribution or temporal stitching can transfer to tactic inference.
-Taken together, these findings underscore the empirical soundness of our benchmark.
+Taken together, these outcomes underscore the empirical soundness of our benchmark.
 Table 2: Ablation study of different frame sampling strategies on Qwen2.5VL-7B. We report the AvgSE, AvgRS and Avg metrics.
@@
 As shown in Table 2, performance with Qwen2.5-VL-7B peaks at 32 frames, while adding more frames yields diminishing or even negative returns, particularly for rally-sequence metrics.
-These results suggest that finer temporal resolution is most beneficial for capturing brief, high-frequency stroke cues, whereas rally level reasoning saturates earlier and is more sensitive to redundancy.
+These results suggest that finer temporal resolution is most beneficial for capturing brief, high-frequency stroke cues, whereas rally-level reasoning saturates earlier and is more sensitive to redundancy.
 Fig. 5: Visualization of performance on three sub-tasks, which require no prior knowledge of tennis, based on the same base model.
@@
   • Enhancing the temporal understanding ability of video reasoning models.
-The experimental results have proved that reasoning can improve rally level understanding, but it has not benefited from fine-grained temporal grounding task.
+The experimental results have proved that reasoning can improve rally-level understanding, but it has not benefited from fine-grained temporal grounding task.
 This suggests that future RL post-training should emphasize temporal understanding and alignment mechanisms rather than relying solely on general video reasoning.